# Detection and Tracking of MAVs Using a Rosette Scanning Pattern LiDAR

### v2

@@ IV-C Losing and regaining Tracking
Figure 6: Particle filter error and $\sigma_{particles}$ are plotted together for a typical loss and regain event. The tracking was lost at around 4 seconds and was regained at 8 seconds.
@@
Fig. 5 shows a typical loss and regain event of the MAV tracking from (a) to (d) while Fig. 6 shows the corresponding error and $\sigma_{particles}$ characteristics. This is one traversal of the drone through the 3 waypoints of the ’lost and found’ pattern. First, the normal MAV tracking is depicted in Fig 5(a). The particles are grouped around the drone with $\sigma_{particles}<0.2\;m$. When the drone starts moving, the error increases slightly but the tracking is still stable. At around 4.5 seconds the drone gets behind the obstacle and $\sigma_{particles}$ increases over $0.2\;m$ meaning the tracking is lost and the particle cloud starts inflating (Fig. 5(b)). Note on the plot, that at first the $\sigma_{particles}$ jumps sharply when the velocity is still estimated and then settles to a much gentler slope when only the prediction noise is added to the particle positions. The particle cloud inflates further to almost $\sigma_{particles}=0.8\;m$ which is depicted in Fig. 5(c) while the error stays constant. This is because the drone hovers behind the obstacle and the estimated pose is the mean of the particle positions which is around the position where the tracking was lost. Then the drone flies out from behind the cover and the particle update and resample steps of the filter condense the cloud instantly which is well depicted on the sharp drop of $\sigma_{particles}$ below $0.2\;m$ in Fig. 6. The tracking is stable again in Fig. 5(d) and the error also decreases to the static error at about 9 seconds when the drone stops moving.
Overall the regaining of tracking after loss is practically instantaneous allowing robust tracking even in scenarios where the trajectory is partially occluded.
V OUTDOOR FIELD TESTS and COMPARISON
After the successful indoor tests, we validated the developed method in an outdoor environment as well. We performed the measurements with the presented method under two different weather conditions (on a sunny and a foggy day). The purpose of the validation is to test the track movement and to determine the detection distances and number of detected points in outdoor conditions, on a commercial drone. Finally, we compared the performance of our method to the previous state-of-the-art [13].
V-A Outdoor test
For outdoor tests, the DJI Phantom drone was used on approximately flat terrain. DJI Phantom 4 Pro V2.0 measures $251\;mm\times 398.78\;mm\times 172.72\;mm$ and weighs $1375\;g$ with a battery and propellers attached. An important parameter for detection is the frame size (the largest distance between the motors), which is $350\;mm$ for this type of drone. The measurements were carried out in the factory state of the drone, we did not improve the detection range by using special paint or other methods. For a video of the outdoor experiment, kindly check the video attachment here: https://youtu.be/VdePm3on8vo.
A sunny day was chosen for the first and a foggy day for the second measurement day when visibility was also severely limited (see Fig.7).
V-B Results
(a) Sunny
(b) Foggy
When evaluating drone detection, an important value is the maximum detection distance. The drone’s trajectory was successfully tracked up to a distance of $128\;m$ in the sunny test environment (Fig. 8(a)), and up to a distance of $58\;m$ in the foggy test environment (Fig. 8(b)). As explained earlier in subsection IV-A, we consider the drone tracked when $\sigma_{particles}<0.2\;m$. We consider the tracking lost after 10 subsequent missing measurements, but it is important to note that we had some occasional detections at distances of up to $145\;m$ on the sunny day.
In both scenarios, we calculated maximum speeds of a little more than $6\;m/s$ from the tracking positions and timestamps.
V-C Comparison
We compare our outdoor test results with another study in which the drone detection efficiency of the VLP-16 LiDAR was tested. In the work of Dogru et al.[13], the size of the drones is larger than the one used for our measurement, which affects the number of reflected points. Furthermore in their paper, there are measurements in which a retroreflector was used to increase the number of points reflected to the sensor.
TABLE I: Comparison of maximum detection distances
Measurement, method Weather With retroreflector Distance($m$)
Our test I. Sunny No 128
Our test II. Foggy No 58
Dogru et al.[13] Sunny Yes 100
Dogru et al.[13] Sunny No 70
Our maximum distance of detection is about $1.5$ to $2$ times greater than their result as in TABLE I.
Another important aspect is the number of detected points, which greatly impacts the robustness of tracking. In sunny weather, except for the last part of the detection range, the number of detected points in the result is $3$ to $5$ times higher than in the current state-of-the-art (see Fig. 9).
Figure 9: Number of detected points depending on distance. The proposed algorithm returns considerably more points in every distance range than [13].
In the case of foggy weather, the detection distance was significantly reduced, as expected in advance based on the literature[39].
VI CONCLUSION
@@
This article presents a novel tracking method utilizing a rosette scanning pattern LiDAR mounted on a Pan-Tilt platform. The proposed algorithm has been rigorously tested in both indoor and outdoor environments, demonstrating accuracy on par with the current state-of-the-art indoor method [24] and considerably surpassing the leading outdoor method [13] in terms of detection range and point density. Notably, our method now tracks all dynamic objects, and ongoing work aims to further enhance the system by differentiating between drones and other flying objects.

### v3

@@ III-F Losing and regaining Tracking
Figure 12: Particle filter error and $\sigma_{particles}$ are plotted together for a typical loss and regain event. The error of the EKF of [13] is also shown for comparison. The tracking was lost at around 2 seconds and was regained at 6 seconds.
@@
FIGURE 11 shows a typical loss and regain event of the MAV tracking from (a) to (d) while FIGURE 12 shows the corresponding error and $\sigma_{particles}$ characteristics. FIGURE 12 also shows the error of the EKF implementation of [13] which varies the NNS radius based on the uncertainty, giving the method re-detection capability. This is one traversal of the drone through the 3 waypoints of the ’Lost and found’ pattern. First, the normal MAV tracking is depicted in FIGURE 11(a). The particles are grouped around the drone with $\sigma_{particles}<\sigma_{threshold}$. When the drone starts moving, the error increases slightly but the tracking is still stable. At around 2 seconds the drone gets behind the obstacle and $\sigma_{particles}$ increases over $\sigma_{threshold}$, meaning the tracking is lost and the particle cloud starts inflating (FIGURE 11(b)). Note on the plot that at first the $\sigma_{particles}$ jumps sharply when the $\rho$ ratio is still considerable and then settles to a much gentler slope when only the prediction noise is added to the particle positions. The particle cloud inflates further to almost $\sigma_{particles}=0.4\;m$, which is depicted in FIGURE 11(c) while the error stays constant. This is because the drone hovers behind the obstacle, and the estimated pose is the mean of the particle positions, which is around the position where the tracking was lost. Then the drone flies out from behind the cover, and the particle update and resample steps of the filter condense the cloud instantly, which is well depicted on the sharp drop of $\sigma_{particles}$ below $\sigma_{threshold}$ in FIGURE 12. The tracking is stable again in FIGURE 11(d), and the error also decreases to the static error at about 6 seconds when the drone stops moving.
Overall, the regaining of tracking after loss is practically instantaneous for our PF, while the EKF recovers slightly slower. This allows robust tracking even when the trajectory is partially occluded. Notably, the PF error is constrained throughout, while the EKF increases constantly until re-detection. This difference is clearly illustrated in TABLE IV, where the error of the EKF of [13] is more than twice that of the delta position PF. The error reported for [24] is less informative in this context, as their EKF implementation lacks re-detection capability. The continuously increasing error of the EKF is not preferred if the turret is moved by the filter prediction, as it may cause the turret to rotate excessively, losing the drone outside the FoV and preventing re-detection altogether.
For a real-world example where re-detection was demonstrated, kindly check out the attached video7.
IV CONCLUSION
@@
This article presents a novel tracking method utilizing a rosette scanning pattern LiDAR mounted on a pan-tilt platform. The proposed algorithm has been rigorously tested and compared to SOTA methods in both indoor and outdoor environments, demonstrating accuracy on par with the current SOTA indoor method [24] and surpassing the leading outdoor method [13] by an $85\;\%$ increase in detection range and at least a twofold improvement in both coverage and returned point counts.
Our method now tracks all dynamic objects, and ongoing work aims to further enhance the system by differentiating between drones and other flying objects. Future research directions also include the exploration and integration of deep learning-based tracking techniques, the analysis of environmental effects like fog or rain on our algorithm, and the development of robust multi-object tracking strategies tailored for the pan-tilt configuration.